OpenStudy (percyjackson&theω):

There are twenty raffle tickets, and four people in a room hold four of those. If two raffle ticket numbers are drawn, without numbers being put back into the pool, what is the percentage of probability that both raffle tickets will belong to people in the room?

OpenStudy (percyjackson&theω):

@matt101

OpenStudy (anonymous):

ok so find the probability it will not happen by doing, 16/20 + 15/19 = ? What do you get?

OpenStudy (anonymous):

i apologize you multiply you dont add that was a typo

OpenStudy (anonymous):

you do 1 - (answer you got) and that should give you the answer

OpenStudy (percyjackson&theω):

Hmm... please keep in mind I'm fairly new to this concept and if you could dumb down your explanations it would be appreciated.... So...

OpenStudy (anonymous):

lol no worries

OpenStudy (percyjackson&theω):

I'm looking for how and why, I have a page of similar questions to do and I'm not sure how to go about doing them..

OpenStudy (anonymous):

ok so 20 people could win we want to know the odds that none of the four are drawn so, the probability one would be drawn at first is 4/20, the second would be 4/19 because it says they do not put the number back

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

should be fairly intuitive...

OpenStudy (percyjackson&theω):

Hmm... I guess so

OpenStudy (anonymous):

it should make sense

OpenStudy (anonymous):

4 / 20 have ticket what are odds one of them get picked 4/20 same for 4/19 what is not clear? happy to explain more

OpenStudy (anonymous):

you there?

OpenStudy (percyjackson&theω):

Why isn't it clear? Maybe because before today I didn't know about this concept... I think... I think I'll try Khan Academy and see what that has to offer.

OpenStudy (anonymous):

alright sorry I couldnt make it more clear

OpenStudy (percyjackson&theω):

No problem, thanks anyway.

OpenStudy (anonymous):

good luck! once you get over first hurdle of probability it all becomes easier :)

OpenStudy (percyjackson&theω):

Ha, I should hope so.

OpenStudy (perl):

does that mean 4 people each hold 4 tickets, so there are 16 tickets held in all by the 4 people?

OpenStudy (percyjackson&theω):

No, I think each of the four people hold one ticket.

OpenStudy (matt101):

Have you figured out the question yet?

OpenStudy (perl):

I would like to verify this, the wording is a bit confusing.

OpenStudy (kropot72):

There are 4C2 combinations of the four tickets taken 2 at a time that are held by the 4 people in the room. The total number of combinations of the 20 tickets in the draw taken 2 at a time is 20C2. Therefore the required probability is given by: \[\large \frac{4C2}{20C2} \times 100=you\ can\ calculate\]